40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 28 17:00:50 2013, Jenne, Update, ASC, Proposed POP path, to be installed this evening CurrentPOPpath.pdfProposedPOPpath.pdf
    Reply  Thu May 30 17:18:58 2013, Jenne, Update, ASC, New POP path 
       Reply  Sat Jun 1 14:14:56 2013, Jenne, Update, ASC, New POP path POX_table_30May2013progress.pdfPOP_QPD_lensSolution.png
          Reply  Fri Jun 14 03:10:40 2013, Jenne, Update, ASC, New POP path - PDs in place, need cabling 
             Reply  Fri Jun 14 17:54:11 2013, Jenne, Update, ASC, New POP path - cabling work 
             Reply  Wed Jun 19 22:02:03 2013, Jenne, Update, ASC, New POP path - ready to try 
Message ID: 8658     Entry time: Thu May 30 17:18:58 2013     In reply to: 8649     Reply to this: 8663
Author: Jenne 
Type: Update 
Category: ASC 
Subject: New POP path 

I have placed the G&H mirrors and the Y1 as pictured in my proposed layout in elog 8649.  The distance between the 2" lens and the PDs has increased, so the focus point is all wrong.  I have measured the distances between optics on the table, and will pick new lenses and finish the POP layout later today or tomorrow.

For now, here are the powers measured using the Ophir power meter:

----------------------

PRM-ITMY lock, POPDC was ~190 counts

5.29 uW after Y1 weird angle.  Can't see beam before then to measure

5.27 uW before BS50
3.5 uW before razor PD
3.00 uW before 110PD

 ------------------------

After installing G&H mirrors, replacing BS-98 with Y1:

4.94 uW before y1 (after G&H's)
4.92 uW after y1
2.66 uW before razor PD
1.61 uW before 110PD

ELOG V3.1.3-